# Bilevel Reinforcement Learning via the Development of Hyper-gradient without Lower-Level Convexity

### v1

@@ Appendix I Details on Experiments
This section presents the details of the experiment implementation. We test on an Atari game, BeamRider, from the Arcade Learning Environment (ALE) [5] to empirically validate the performance of the model-free algorithm, SoBiRL. The reward provided by ALE serves as the ground truth, and for each trajectory pair, preference is assigned to the trajectory with higher accumulated ground-truth reward. We compare SoBiRL with DRLHF [14] and a baseline algorithm SAC [29, 30]. Both SoBiRL and DRLHF harness deep neural networks to predict rewards, while the baseline SAC receives ground-truth rewards for training. The experiments are produced on a server that consists of two Intel® Xeon® Gold 6330 CPUs (total 2$\times$28 cores), 512GB RAM, and one NVIDIA A800 (80GB memory) GPU.
I.1 Practical SoBiRL
@@
SoBiRL adopts deep neural networks to parameterize the policy and the reward model. In this way, we supplement Algorithm 2 with details, resulting in the practical version, Algorithm 3. The SAC algorithm capable of automating the temperature parameter [30] is chosen as the lower-level solver for SoBiRL, i.e., for a fixed reward network, SAC is invoked with several timesteps to update the policy network, the Q-value network and the adaptive temperature parameter (line 2 of Algorithm 3). Here, to focus more on the implementation of SoBiRL, we stepsize the details of SAC, which fully follows the principles in [30].
@@
3:     Rollout the policy $\pi_{k}=\pi_{\theta^{\pi}_{k}}$ in the uppper-level MDP $\triangleright$ collect trajectories
4:     to obtain trajectory pairs
5:     Query preference and update the buffer $\mathcal{D}$ with the labeled pairs
6:     Compute the estimator ¯$\overline{\nabla}\phi\left(\theta^{r}_{k},\pi_{k},\tau_{k}\right)$ using $\mathcal{D}$triangleright upper-level reward update
7:     Implement an inexact hyper-gradient descent step
8:     $\theta^{r}_{k+1}=\theta^{r}_{k}-\beta\overline{\nabla}\phi\left(\theta^{r}_{k},\pi_{k},\tau_{k}\right)$
9:  end for
9:  $(\theta^{r}_{K+1},\theta^{\pi}_{K+1})$
Considering the hyper-gradient estimator, we also introduce a practical version, ¯$\overline{\nabla}\phi\left(x_{k},\pi_{k},\tau_{k}\right)$, adapted from ~$\widetilde{\nabla}\phi\left({x_{k},\pi_{k}}\right)$ in Section 5.
¯$\overline{\nabla}\phi\left(x_{k},\pi_{k},\tau_{k}\right)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{k}\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x_{k}\right)\right]$ (78)
$+\tau_{k}^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{k}\right)}\left[l\left(d_{1},d_{2},...,d_{I};x_{k}\right)\left(\sum_{i}{\sum_{h}{\nabla\left(r_{s_{h}^{i}a_{h}^{i}}(x_{k})-\mathbb{E}_{a^{\prime}\sim\pi_{k}\left(\cdot|s_{h}^{i}\right)}\left[r_{s_{h}^{i}a^{\prime}}(x_{k})\right]\right)}}\right)\right].$
The main difference between (78) and (15) lies in two aspects: 1) ¯$\overline{\nabla}\phi\left(x_{k},\pi_{k},\tau_{k}\right)$ accommodates the adaptive temperature parameter $\tau_{k}$; 2) the term $\nabla\left(r_{s_{h}^{i}a_{h}^{i}}(x_{k})-\mathbb{E}_{a^{\prime}\sim\pi_{k}\left(\cdot|s_{h}^{i}\right)}\left[r_{s_{h}^{i}a^{\prime}}(x_{k})\right]\right)$ is employed as an one-depth truncation of ~$\widetilde{\nabla}\left(Q_{s_{h}^{i}a_{h}^{i}}\left({x_{k},\pi_{k}}\right)-V_{s_{h}^{i}}\left({x_{k},\pi_{k}}\right)\right)$ in (15), inspired by the expressions,
@@
I.2 Experiment settings
Wrappers are employed on the Atari game, which originate from [60]: initial 00 to $30$ no-ops to inject stochasticity, max-pooling pixel values over the last two frames, an episodic life counter, four-frame skipping to accelerate sampling, four-frame stacking to help infer game dynamics, warping the image to size $84\times 84$ and clipping rewards to $[-1,1]$.
As mentioned above, we take SAC as the lower-level solver for both SoBiRL and DRLHF. All the actor network, the critic network, and the temperature parameter are updated by the Adam optimizer, with an initial learning rate $10^{-4}$ linearly decaying to 00 after $8\times 10^{6}$ time steps (although the runs were actually trained for only $4\times 10^{6}$ timesteps). In addition, we set the initial temperature parameter $\tau_{0}=1$, the inner iteration timesteps $N=10^{4}$, and the batch size to $64$.
Both SoBiRL and DRLHF employ deep neural networks to parameterize the reward model. The wrapped environment returns an $84\times 84\times 4$ tensor as the state information. Therefore, data in the shape of $84\times 84\times 4$ is fed into the reward network. It undergoes four convolutional layers with kernel sizes of 7×7, 5×5, 3×3, 3×3775533337$\times$ 7,\ 5$\times$ 5,\ 3$\times$ 3,\ 3$\times$ 37 × 7 , 5 × 5 , 3 × 3 , 3 × 3 and stride values of 3, 2, 1, 132113,\ 2,\ 1,\ 13 , 2 , 1 , 1, respectively. Each convolutional layer holds $16$ filters and incorporates leaky ReLU nonlinearities ($\alpha=0.01$). Subsequently, the data passes through a fully connected layer of size $64$ and is then transformed into a scalar. Batch normalization and dropout with a dropout rate of $0.5$ are applied to all convolutional layers to mitigate overfitting. An AdamW [55] optimizer is adopted with the learning rate $3\times 10^{-4}$, betas $(0.9,0.999)$, epsilon $10^{-8}$ and weight decay $10^{-2}$.

### v2

@@ I.3 Convergence Analysis of Stoc-SoBiRL
This section presents the details of the experiment implementation. We conduct the RLHF task on three Atari games from the Arcade Learning Environment (ALE) (Bellemare et al.,, 2013) and three Mujoco environments (Todorov et al.,, 2012) to empirically validate the performance of the model-free algorithm, SoBiRL. The reward provided by the original environments serves as the ground truth, and for each trajectory pair, preference is assigned to the trajectory with higher accumulated ground-truth reward. We compare SoBiRL with the bilevel algorithms DRLHF (Christiano et al.,, 2017), PBRL (Shen et al.,, 2024), HPGD Thoma et al., (2024), and a baseline algorithm SAC (Haarnoja et al., 2018a, ; Haarnoja et al., 2018b, ). All the bilevel solvers harness deep neural networks to predict rewards, while the baseline SAC receives ground-truth rewards for training. Additionally, we also test on a synthetic bilevel RL experiment to verify the convergence rate of the model-based algorithm, M-SoBiRL. The experiments are produced on a workstation that consists of two Intel® Xeon® Gold 6330 CPUs (total 2$\times$28 cores), 512GB RAM, and one NVIDIA A800 (80GB memory) GPU. We have made the code available on https://github.com/UCAS-YanYang/SoBiRL.
J.1 Practical SoBiRL
@@
SoBiRL adopts deep neural networks to parameterize the policy and the reward model. In this way, we supplement Algorithm 1 with details, resulting in the practical version, Algorithm 5. The SAC algorithm capable of automating the temperature parameter (Haarnoja et al., 2018b, ) is chosen as the lower-level solver for SoBiRL, i.e., for a fixed reward network, SAC is invoked with several timesteps to update the policy network, the Q-value network and the adaptive temperature parameter (line 2 of Algorithm 5). Here, to focus more on the implementation of SoBiRL, we stepsize the details of SAC, which fully follows the principles in (Haarnoja et al., 2018b, ).
@@
3:     Rollout the policy $\pi_{k}=\pi_{\theta^{\pi}_{k}}$ in the uppper-level MDP $\triangleright$ collect trajectories to obtain trajectory pairs
4:     Query preference and update the buffer $\mathcal{D}$ with the labeled pairs
5:     Compute the estimator $\nabla_{p}\phi\left(\theta^{r}_{k},\pi_{k},\tau_{k}\right)$ using $\mathcal{D}$triangleright upper-level reward update
6:     Implement an inexact hyper-gradient descent step $\theta^{r}_{k+1}=\theta^{r}_{k}-\beta\nabla_{p}\phi\left(\theta^{r}_{k},\pi_{k},\tau_{k}\right)$
7:  end for
7:  $(\theta^{r}_{K+1},\theta^{\pi}_{K+1})$
Adapting from ~$\widetilde{\nabla}\phi\left({x_{k},\pi_{k}}\right)$ in Section 5, a practical version of hyper-gradient estimator $\nabla_{p}\phi\left(x_{k},\pi_{k},\tau_{k}\right)$ is used,
$\nabla_{p}\phi\left(x_{k},\pi_{k},\tau_{k}\right)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{k}\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x_{k}\right)\right]$ (99)
$+\tau_{k}^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{k}\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x_{k}\right)\left(\sum_{i}{\sum_{h}{\nabla\left(r_{s_{h}^{i}a_{h}^{i}}(x_{k})-\mathbb{E}_{a^{\prime}\sim\pi_{k}\left(\cdot|s_{h}^{i}\right)}\left[r_{s_{h}^{i}a^{\prime}}(x_{k})\right]\right)}}\right)\right].$
The main difference between (99) and (14) lies in two aspects: 1) $\nabla_{p}\phi\left(x_{k},\pi_{k},\tau_{k}\right)$ accommodates the adaptive temperature parameter $\tau_{k}$; 2) the term $\nabla\left(r_{s_{h}^{i}a_{h}^{i}}(x_{k})-\mathbb{E}_{a^{\prime}\sim\pi_{k}\left(\cdot|s_{h}^{i}\right)}\left[r_{s_{h}^{i}a^{\prime}}(x_{k})\right]\right)$ is employed as an one-depth truncation of ~$\widetilde{\nabla}\left(Q_{s_{h}^{i}a_{h}^{i}}\left({x_{k},\pi_{k}}\right)-V_{s_{h}^{i}}\left({x_{k},\pi_{k}}\right)\right)$ in (14), inspired by the expressions,
@@
J.2 Experiment Settings for RLHF
The results on the Atari games—BeamRider, Seaquest, and SpaceInvaders—are shown in Figures 1 and 2. The results on the Mujoco environments—HalfCheetah, Walker2d, and Hopper—are presented in Figure 3. The overall performance is summarized in Table 2.
Figure 2: Comparison on the Atari games—Seaquest and SpaceInvaders—evaluated by the ground-truth reward. Each bilevel algorithm collects a total of $3000$ trajectory pairs. The results are averaged over $5$ seeds.
Figure 3: Comparison of algorithms on the Mujoco simulations—HalfCheetah, Walker2d, and Hopper—evaluated by the ground-truth reward. Each bilevel algorithm collects a total of $3000$ trajectory pairs. The results are averaged over $5$ seeds.
As mentioned above, we take SAC as the lower-level solver for the bilevel algorithms. We set the initial temperature parameter $\tau_{0}=1$, the inner iteration timesteps $N=10^{4}$ for Atari games and $N=5\times 10^{3}$ for Mujoco simulations, and the batch size to $64$. In addition, on Atari games, the actor network, the critic network, and the temperature parameter are updated by the Adam optimizer, with an initial learning rate $1\times 10^{-4}$ for BeamRider and SpaceInvaders, and $3\times 10^{-4}$ for Seaquest, linearly decaying to 00 after $8\times 10^{6}$ time steps (although the runs were actually trained for only $4\times 10^{6}$ timesteps). On all Mujoco simulations, the critic network and the temperature parameter share the same initial learning rate $1\times 10^{-3}$, while that of the actor network is $3\times 10^{-4}$. They are updated by Adam optimizers, linearly decaying the learning rates to 00 after $8\times 10^{6}$ time steps (although the runs were actually trained for only $4\times 10^{6}$ timesteps).
Bilevel algorithms employ deep neural networks to parameterize the reward model. The wrapped environment returns an $84\times 84\times 4$ tensor as the state information. Therefore, data in the shape of $84\times 84\times 4$ is fed into the reward network. It undergoes four convolutional layers with kernel sizes of 7×7, 5×5, 3×3, 3×3775533337$\times$ 7,\ 5$\times$ 5,\ 3$\times$ 3,\ 3$\times$ 37 × 7 , 5 × 5 , 3 × 3 , 3 × 3 and stride values of 3, 2, 1, 132113,\ 2,\ 1,\ 13 , 2 , 1 , 1, respectively. Each convolutional layer holds $16$ filters and incorporates leaky ReLU nonlinearities ($\alpha=0.01$). Subsequently, the data passes through a fully connected layer of size $64$ and is then transformed into a scalar. Batch normalization and dropout with a dropout rate of $0.5$ are applied to all convolutional layers to mitigate overfitting. An AdamW (Loshchilov and Hutter,, 2018) optimizer is adopted with the learning rate $3\times 10^{-4}$, betas $(0.9,0.999)$, epsilon $10^{-8}$ and weight decay $10^{-2}$.
@@
Regarding the Atari games, wrappers are employed, which originate from (Mnih et al.,, 2015): initial 00 to $30$ no-ops to inject stochasticity, max-pooling pixel values over the last two frames, an episodic life counter, four-frame skipping to accelerate sampling, four-frame stacking to help infer game dynamics, warping the image to size $84\times 84$ and clipping rewards to $[-1,1]$.
Specifically, after conducting initial tests, we observe that PBRL with the “value penalty" and the penalty parameter as $2$ provide robust performance, for which it is adopted for comparison.
J.3 Experiment Settings for the Synthetic Bilevel RL Problem
To test M-SoBiRL, we build a synthetic bilevel RL problem in the form of (4) with the upper-level $f$ as a quadratic function of $(x,\pi)$ and the dimension of $x$ as $n=100$. The lower-level problem is adapted from RL problems used in (Zhan et al.,, 2023) and (Li et al., 2024b, ). Specifically, we take $|\mathcal{S}|=10,|\mathcal{A}|=5$ and $\gamma=0.5,\tau=1$. The transition probabilities are generated randomly, and for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, the reward $r_{sa}$ is parameterized by a quadratic form of $x$ with a random perturbation. The step sizes in Algorithm 2 are taken as $(\beta,\eta)=(0.008,0.5)$. Results of the synthetic experiment are shown in Figure 4. The curves exhibit the benign convergence property of the proposed algorithm.
Figure 4: A synthetic bilevel RL problem to verify the model-based algorithm, M-SoBiRL. Metrics are the hyper-gradient norm $\left\|{\nabla\phi(x)}\right\|_{2}$ and the upper-level loss $f(x,\pi)$.
Appendix K APPLICATIONS OF BILEVEL REINFORCEMENT LEARNING
In this section, we provide a detailed explanation of how the formulation (4), with the upper-level function
$f\left(x,\pi\right)=\mathbb{E}_{d_{i}\sim\rho\left(d;\pi\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right],$ (100)
incorporates bilevel RL applications, using additional examples similar to those in Section 3.3. We also explain how the boundedness of $l$ assumed in Section 7 can be verified individually for each case.
Reward Shaping From the perspective of bilevel RL, we can shape an auxiliary reward function $r(x)$ at the lower level for efficient agent training, while maintaining the original environment at the upper level to align with the initial task evaluation (Hu et al.,, 2020), which is established as follows,
$\min_{x\in\mathbb{R}^{n}}$ $\ -V_{\bar{\mathcal{M}}_{\bar{\tau}}}^{\pi^{*}\left(x\right)}\left(\bar{\rho}\right)=-\mathbb{E}\left[{Q^{\pi^{*}(x)}_{\bar{\mathcal{M}}_{\bar{\tau}}}(s,a)\,|\,s\sim\bar{\rho},\pi^{*}(x)}\right]$
$\mathrm{s.\,t.}$ $\ \ \pi^{*}\left(x\right)=\operatorname*{arg\,min}_{\pi\in\Delta^{\left|{\mathcal{A}}\right|}}-V_{\mathcal{M}_{\tau}\left(x\right)}^{\pi}\left(\rho\right).$
Note that its upper-level function falls into the formulation (100) with $I=1$, $H=1$, $l=-Q^{\pi}_{\bar{\mathcal{M}}_{\bar{\tau}}}\left({s_{0}^{1},a_{0}^{1}}\right)$. Therefore, the boundedness and smoothness of $l$ can be seen from Proposition H.5.
Reinforcement Learning from Human Feedback The target of RLHF is to learn the intrinsic reward function that incorporates expert knowledge, from simple labels only containing human preferences. It optimizes a policy under $r(x)$ at the lower level, and adjusts $x$ to align the preference predicted by the reward model $r(x)$ with the true labels at the upper level.
$\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{n}}&\mathbb{E}_{y,d_{1},d_{2}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l_{h}\left(d_{1},d_{2},y;x\right)\right]\\ \mathrm{s.\,t.}&\pi^{*}\left(x\right)=\operatorname*{arg\,min}\limits_{\pi\in\Delta^{\left|{\mathcal{A}}\right|}}-V_{\mathcal{M}_{\tau}\left(x\right)}^{\pi}\left(\rho\right),\end{array}$
$P(d_{i})=\bar{\rho}\left({s_{0}^{i}}\right)\left[{\Pi_{h=0}^{H-2}\pi^{*}(x)\left({a_{h}^{i}|s_{h}^{i}}\right)\bar{P}\left({s_{h+1}^{i}|s_{h}^{i},a_{h}^{i}}\right)}\right]\times\pi^{*}(x)\left({a_{H-1}^{i}|s_{H-1}^{i}}\right),$
and the preference label $y\in\{0,1\}$, indicating preference for $d_{1}$ over $d_{2}$, obeys human feedback distribution $y\sim D_{human}(y|d_{1},d_{2})$. Moreover, $l_{h}$ is the binary cross-entropy loss, $l_{h}\left(d_{1},d_{2},y;x\right)=-y\log P\left(d_{1}\succ d_{2};x\right)-\left(1-y\right)\log P\left(d_{2}\succ d_{1};x\right)$, with the preference probability $P\left({d_{1}\succ d_{2};x}\right)$ built by the Bradley–Terry model,
$P\left(d_{1}\succ d_{2};x\right)=\frac{\exp\left(\sum_{h=0}^{H-1}{r_{s_{h}^{1},a_{h}^{1}}\left(x\right)}\right)}{\exp\left(\sum_{h=0}^{H-1}{r_{s_{h}^{1},a_{h}^{1}}\left(x\right)}\right)+\exp\left(\sum_{h=0}^{H-1}{r_{s_{h}^{2},a_{h}^{2}}\left(x\right)}\right)}.$ (101)
The upper-level function of RLHF is an instance of (100) with $I=2$, finite $H$, and $l=\mathbb{E}_{y}\left[{l_{h}\left({d_{1},d_{2},y;x}\right)}\right]$. Moreover, the boundedness of the reward $\left|{r_{sa}(x)}\right|\leq C_{r}$ will guarantee the boundedness of $P$, i.e., $\frac{1}{2}{\exp(-2HC_{r})}\leq{P(d_{1}\succ d_{2};x)}\leq 1$, and thus the boundedness of $l$.
Contract Design This problem involves designing payment mechanisms that enable a principal to influence the agent’s decision-making process (Wu et al.,, 2024). Through the state-contingent payment $x\in\mathbb{R}^{\left|{\mathcal{S}}\right|\times\left|{\mathcal{S}}\right|}_{+}$, the principal seeks to incentivize the agent to adopt policies that serve the principal’s interests. Specifically, when the agent takes action $a$ at state $s$, it costs $c(s,a)$ in the lower level while the principle only observes the transition $s\to s^{\prime}$ and receives the reward $R(s,s^{\prime})$. To encourage the transition, the principal offers a positive payment $x(s,s^{\prime})$, which is thus added to the lower-level objective and subtracted from the upper-level objective.
$\min_{x\in\mathbb{R}^{n}}$ $\ -\mathbb{E}\left[{\sum_{t=0}^{H}R(s_{t},s_{t+1})-x(s_{t},s_{t+1})\ |\ s_{0}\sim\rho,\pi^{*}(x)}\right]$
$\mathrm{s.\,t.}$ $\ \ \pi^{*}\left(x\right)=\operatorname*{arg\,min}_{\pi\in\Delta^{\left|{\mathcal{A}}\right|}}-\mathbb{E}\left[{\sum_{t=0}^{\infty}\gamma^{t}(x(s_{t},s_{t+1})-c(s_{t},a_{t})+\tau h(\pi_{s_{t}}))\ |\ s_{0}\sim\rho,\pi}\right].$
Note that its upper-level function is an instance of (100) with $I=1$, finite $H$, and $l=\sum_{t=0}^{H}-R(s_{t},s_{t+1})+x(s_{t},s_{t+1})$. It is reasonable that the reward $R$ and the payment $x$ are bounded in practice, and thus the function $l$ is bounded.
Efficient Robot Navigation As formulated in (Chakraborty et al.,, 2024), we consider a maze-world environment of size $N\times N$, and aim to navigate a robot to get close to the destination $(N,N)$ efficiently. To this end, the designer can introduce a modified goal $x\in\mathbb{R}^{2}$, which induces an associated reward $r_{x}\in\mathbb{R}^{\left|{\mathcal{S}}\right|\times\left|{\mathcal{A}}\right|}$, and then use this goal to train the robot (in the lower level). In this manner, the upper-level objective includes both the received reward and the additional cost of moving the robot from $x$ to $(N,N)$. The task is formulated by the following bilevel problem,
$\min_{x\in\mathbb{R}^{2}}$ $\ \omega\left\|{x-(N,N)}\right\|^{2}-\mathbb{E}\left[{\sum_{t=0}^{H}r_{x}(s_{t},a_{t})|\ s_{0}\sim\rho,\pi^{*}(x)}\right]$
$\mathrm{s.\,t.}$ $\ \ \pi^{*}\left(x\right)=\operatorname*{arg\,min}_{\pi\in\Delta^{\left|{\mathcal{A}}\right|}}-\mathbb{E}\left[{\sum_{t=0}^{\infty}\gamma^{t}(r_{x}(s_{t},a_{t})+\tau h(\pi_{s_{t}}))\ |\ s_{0}\sim\rho,\pi}\right],$
where $\omega>0$ is the weight balancing the deviation from the original destination $(N,N)$ and the accumulative reward measured under the modified goal $x$. In this case, the upper-level function matches (100) with $I=1$, finite $H$, and $l=-\sum_{t=0}^{H}\left({r_{x}(s_{t},a_{t})-\omega\left\|{x-(N,N)}\right\|^{2}/H}\right)$, which is bounded if the reward is bounded.
Appendix L EXTENSION TO ASYNCHRONOUS AND DISTRIBUTED BILEVEL RL SETTINGS
In Algorithm 1, the task within the $k$-th outer iteration is divided into two components: one is dedicated to approximately solving the lower-level problem such that $\|\pi_{k}^{*}-\pi_{k}\|^{2}_{2}\leq\epsilon$, and the other is used to collect rollouts via $\pi_{k}$. We denote the sample complexity as $c_{\mathrm{solver}}$ and $c_{\mathrm{roll}}$, respectively. Employing the asynchronous strategy, existing work (Shen et al.,, 2023) shows the actor-critic algorithm achieves a sample complexity $c_{\mathrm{solver}}/{N}$ per worker, where $N$ is the number of workers. In a similar spirit, we can also take advantage of multiprocessing to collect rollouts amongst $N$ workers asynchronously and thus enjoy $c_{\mathrm{roll}}/N$ complexity per worker. Finally, aggregate the rollouts from all the workers to estimate the hyper-gradient and update the upper-level variable. In this manner, the error bound for the hyper-gradient estimation in our work still holds, and thus the convergence analysis can be generalized.
In the distributed setting (e.g., see (Chen et al., 2022c, )), agents are connected via a fully decentralized network and interact with a shared environment. Since the agent can communicate only with its neighbors, estimating the (global) hyper-gradient becomes challenging. Therefore, developing a tailored communication strategy and establishing a convergence analysis is an interesting and meaningful topic with broad applications in multi-agent RL problems. To this end, recent advances in decentralized bilevel optimization offer valuable insights. (Kong et al.,, 2024; He et al.,, 2024; Zhu et al.,, 2024).